OpenStudy (anonymous):

ln(x + 3) − ln(x) = 3. solve analytically??

OpenStudy (anonymous):

that was wrong! start with \[\log(\frac{x+3}{x})=3\] then equivalent exponential form gives \[\frac{x+3}{x}=10^3\] so \[\frac{x+3}{x}=1000\] \[x+3=1000x\] etc

OpenStudy (anonymous):

ln [(x + 3)/x] = 3 ===> e^3 = [(x + 3)/x] ===> e^3 (x) =x+3 ==> e^3 (x) -x=3 ==> x(e^3-1)=3 ===> x=3/(e^3-1)

OpenStudy (anonymous):

The problem is written as a natural log, so raheen's answer is correct.

OpenStudy (anonymous):

ah so it is.

OpenStudy (anonymous):

raheen is right

OpenStudy (anonymous):

yes raheen is right for sure